If 2=10^b, write down in terms of b, expressions for
a) 4
b) 16
c) 0.25

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

2=10^b
log 2=b ANSWER a
log 4=log 2^2=2 log 2. 10^2 log 2=4, but log 2=b, so 10^2b=4 ANSWER b
16=4b, log 16=4b log 10, so 16=4b
0.25=10^b
(1/4)=10^b
log(1/4)=-log 4
-log 4 is -log 2^2=-2 log 2 which is -2b, so 10^-2b=0.25 ANSWER c

Answer by greenestamps(13215)